# Nicolas Crampé1, Rafael I. Nepomechie2, Luc Vinet3,4,5

### v2

@@
Nicolas Crampé1, Rafael I. Nepomechie2, Luc Vinet3,4,5
and Nabi Zare Harofteh2
1CNRS – Université de Montréal CRM - CNRS, France
2Department of Physics, PO Box 248046
@@
3Centre de Recherches Mathématiques, Université de Montréal
@@
4Département de Physique, Université de Montréal
@@
5IVADO, 6666 Rue Saint-Urbain
@@ 4 Discussion
We have seen that, after suitably adjusting a uniform transverse magnetic field, the generic inhomogeneous open XX spin chain has a two-fold degeneracy, and an enhanced $su(2)$ symmetry (1.3). The “inhomogeneous” symmetry generators (1.4) depend on coefficients $a_{n}$ (2.15) that can be computed explicitly for models associated with discrete orthogonal polynomials. The expressions for the generators (1.4), with their non-local “tails” of $\sigma^{z}$ operators, are reminiscent of coproducts of quantum-group generators [11]. The symmetry uncovered here may provide further insight into quench dynamics [7, 8] and protected phases [12] in inhomogeneous spin chains. It would be very interesting if this type of symmetry could be extended to XXZ-type chains. It would also be interesting to find higher-rank algebras with such “inhomogeneous” generators, and corresponding physical models with such symmetries.
@@ Acknowledgements
NC is supported in part by the international research project AAPT of the CNRS. RN is supported in part by the National Science Foundation under grant PHY 2310594, and by a Cooper fellowship. LV is funded in part by a Discovery Grant from the Natural Sciences and Engineering Research Council (NSERC) of Canada.

### v3

# and Nabi Zare Haroftehb

@@
Nicolas Crampéa, Rafael I. Nepomechieb, Luc Vinetc,d,e 111email:luc.vinet@umontreal.ca (Corresponding author)
and Nabi Zare Haroftehb
aCNRS – Université de Montréal CRM - CNRS, France
bDepartment of Physics, PO Box 248046
@@
cCentre de Recherches Mathématiques, Université de Montréal
@@
dDépartement de Physique, Université de Montréal
@@
eIVADO, 6666 Rue Saint-Urbain
@@ 4 Discussion
We have seen that, after suitably adjusting a uniform transverse magnetic field, the generic inhomogeneous open XX spin chain has a two-fold degeneracy, and an enhanced $su(2)$ symmetry (1.3). The “inhomogeneous” symmetry generators (1.4) depend on coefficients $a_{n}$ (2.15) that can be computed explicitly for models associated with discrete orthogonal polynomials. The expressions for the generators (1.4), with their non-local “tails” of $\sigma^{z}$ operators, are reminiscent of coproducts of quantum-group generators [11]. The exact two-fold degeneracy of all the levels of the open XX chain due to the $su(2)$ symmetry presented here is reminiscent of the approximate pairing of all the levels in the ordered phase of the one-dimensional open Ising chain due to the presence of a strong zero mode [12, 13, 14]. The edge spins of the latter model are known to have long coherence times [14]. For the former model, perhaps some (presumably nonlocal) observable could be found, which has suitable overlap with the $su(2)$ charges and is therefore conserved for long times. The symmetry uncovered here may provide further insight into quench dynamics [7, 8] and protected phases [15] in inhomogeneous spin chains. It would be very interesting if this type of symmetry could be extended to XXZ-type chains. It would also be interesting to find higher-rank algebras with such “inhomogeneous” generators, and corresponding physical models with such symmetries.
@@ Acknowledgements
We thank Eric Vernier for helpful correspondence. NC is supported in part by the international research project AAPT of the CNRS. RN is supported in part by the National Science Foundation under grant PHY 2310594, and by a Cooper fellowship. LV is funded in part by a Discovery Grant from the Natural Sciences and Engineering Research Council (NSERC) of Canada.
Conflict of interest
On behalf of all authors, the corresponding author states that there is no conflict of interest.
Data availability
This manuscript has no associated data.